OpenStudy (anonymous):

Two legs of a triangle measure 7 m and 13 m. How many meters long is the hypotenuse? Answer 20 6

Parth (parthkohli):

The Pythagorean!

Parth (parthkohli):

\[a^2 + b^2 = c^2 \Longrightarrow 7^2 + 13^2 = c^2 \Longrightarrow c = \sqrt{7^2 + 13^2} \]

Parth (parthkohli):

Get my point there?

OpenStudy (anonymous):

ya